OpenStudy (highschoolmom2010):

Find the value of the variable. If your answer is not an integer, leave it in simplest radical form.

OpenStudy (highschoolmom2010):



OpenStudy (highschoolmom2010):

\[\frac{ \sqrt[2]{3} }{ 2 }\] \[\sqrt[5]{3}\] \[\sqrt[5]{2}\] \[\frac{ \sqrt[5]{2} }{ 2 }\]

OpenStudy (shamil98):

45-45-90 special right triangles. do you remember the rules and such for it?

OpenStudy (shamil98):

h = L√2

OpenStudy (highschoolmom2010):

kinda

OpenStudy (shamil98):

5 = L√2 5/√2 = L now remember you cant have a number square rooted as the demoninator so 5/√2 x √2/√2 = 5 √2/2

OpenStudy (highschoolmom2010):

\[\frac{ 5 \sqrt2 }{ 2}\]

OpenStudy (shamil98):

Yes.